If the gases in our atmosphere (nitrogen, oxygen, carbon dioxide and water vapor) were lined up to run a race, who would come in 1st?  Which one would be last?  Explain using Graham's Law. Remember that some of those gases are diatomic.


1
Expert's answer
2021-03-03T01:59:41-0500

According to Graham's law, the rate of effusion or diffusion of a gas is inversely proportional to the square root of its molar mass:

"\frac{R_1}{R_2}=\sqrt{\frac{M_2}{M_1}}"

Therefore, the greater the molar mass of a gas, the slower it is going to "run a race". Their molar masses are:

N2 - 28.01 g/mol

O2 - 32.00 g/mol

CO2 - 44.01 g/mol

H2O - 18.02 g/mol

Therefore, the water vapor would be the 1st, and carbon dioxide would be the last.